# COBias and Debias: Minimizing Language Model Pairwise Accuracy Bias via Nonlinear Integer Programming

### v1

@@ Abstract
For language model classification, would you prefer having only one workable class or having every class working? The latter makes more practical uses. Especially for large language models (LLMs), the fact that they achieve a fair overall accuracy by in-context learning (ICL) obscures a large difference in individual class accuracies. In this work, we uncover and tackle language models’ imbalance in per-class prediction accuracy by reconceptualizing it as the Contextual Oddity Bias (COBias), and we are the first to engage nonlinear integer programming (NIP) to debias it. Briefly, COBias refers to the difference in accuracy by a class A compared to its “odd” class, which holds the majority wrong predictions of class A. With the COBias metric, we reveal that LLMs of varied scales and families exhibit large per-class accuracy differences. Then we propose Debiasing as Nonlinear Integer Programming (DNIP) to correct ICL per-class probabilities for lower bias and higher overall accuracy. Our optimization objective is directly based on the evaluation scores by COBias and accuracy metrics, solved by simulated annealing. Evaluations on three LLMs across seven NLP classification tasks show that DNIP simultaneously achieves significant COBias reduction (-27%) and accuracy improvement (+12%) over the conventional ICL approach, suggesting that modeling pairwise class accuracy differences is a direction in pushing forward more accurate, more reliable LLM predictions.
@@ 1 Introduction
We rethink language models’ imbalance in per-class prediction accuracy and reconceptualize it as the Contextual Oddity Bias (COBias). In a nutshell, COBias refers to the difference in accuracy by a class A compared to its “odd” class, which holds the majority wrong predictions of class A. COBias reflects a common language model failure, seen in a variety of NLP benchmark tasks from text classification to question answering. Especially for large language models (LLMs), the fact that they achieve remarkable test accuracy over all class labels by few-shot in-context learning (ICL) [1] can obscure a large difference in individual class accuracies. It is a direct outcome of LLMs’ liking of specific patterns or answers, leading to highly imbalanced predicted class distributions [2, 3]. This per-class accuracy difference can be further attributed to ICL prompt sensitivity or rooted in the black-box pretraining data. To name a few, the causes encompass predilections for common tokens in the pretraining corpora [4], choices of prompt templates [5, 6, 7], demonstrations [8], and prompt orders [9, 10]. Unintended LLM behaviors may also exacerbate the bias, such as user-following sycophancy [11, 12, 13] and prompt toning effects [14, 15, 16, 17]. Despite the causes, COBias renders a language model less effective for lower-accuracy classes and hinders users from trusting the answers, heightening the need for a unified debiasing method at output level.
Importantly, it is not just that a model tends to over-predict a class, but more precisely, the model tends to predict certain classes (not always all others) as it. That is why we need to model the interaction between pairwise class predictions for better bias reduction. An example of an odd class and the per-class accuracy difference are shown by the upper plots in Figure 1. Essentially, mitigating per-class accuracy bias is a critical step in pushing forward LLM inference abilities, as it directly targets improving lower-accuracy classes while improving or at least not hurting the overall accuracy.
In this work, we first put forward COBias and propose the COBias evaluation metric, which measures the averaged pairwise class accuracy difference (Section 2). On three widely used, different-scale LLMs including GPT-2-XL [18], Llama-2-7B and Llama-2-13B [19], we manifest that large COBias consistently exists across seven varied NLP classification tasks.
To reduce COBias and rectify LLM outputs, we directly operate on the per-class probabilities given by the conventional ICL approach. We propose DNIP, the Debiasing as Nonlinear Integer Programming method, to adjust per-class probabilities towards lower COBias and higher overall accuracy, from a combinatorial optimization viewpoint (Section 3). We formulate the debiasing problem as a discrete correction weight selection problem. In particular, DNIP re-weights per-class probability with an optimal set of weights selected based on the evaluation metrics of accuracy and COBias, and thus obtains much less biased predictions with higher accuracy.
On three LLMs across seven NLP evaluation datasets, we show the effectiveness of DNIP. Our key key messages are as follows:
  • We introduce the COBias evaluation metric to assess the pairwise class accuracy bias in language model predictions, and reveal that LLMs of different scales and families consistently exhibit a large COBias score, at an average of 43%.
@@
  • Experiments show that DNIP achieves the best of both worlds for lower COBias (avg. $43\%\rightarrow 16\%$) and higher overall accuracy (avg. $52\%\rightarrow 64\%$) over the ICL approach.
2 The Contextual Oddity Bias and the Debiasing Problem
@@
We conceptualize an odd class to be a class at inference time where predictions of another class are often biased towards it. An odd class is relative to the other class. The existence of odd classes manifests the per-class accuracy bias. We term the accuracy difference between a class A and its odd class as the Contextual Oddity Bias (COBias), suggesting a model’s tendency to over-predict an answer that is caught as odd than the correct answer. Notably, an odd answer may reflect surface patterns captured from the task-specific prompts or a common token in the pretraining data; therefore, even on the same task, which answers are odd can vary across different models. That’s why we use “contextual” to mean that the oddity of an answer depends on the model or task. In addition, the prefix “CO” in COBias can be interpreted as the pairwise nature of the bias, suggesting that if there is an odd class, there has to be a class that is biased to it.
@@
The Contextual Oddity Bias Metric. We define the contextual oddity bias metric as follows. Formally, given an $N$-way classification task, let $x_{m}$ denote an input whose label is $y_{m}$. By few-shot prompting, let $\boldsymbol{\boldsymbol{p}_{m}}=(p_{m1},\dots,p_{mN})$ denote the output token probability over $N$ classes $c_{1},\dots,c_{N}$. The ICL prediction $\hat{y}_{m}$ is $\operatorname*{arg\,max}_{j\in\{1,\dots,N\}}\boldsymbol{p}_{m}$. For a given class $c_{i}$, we denote the class it is most biased towards as its oddest class $c_{o}$. Suppose test accuracy for the two is $A_{o}$ and $A_{i}$ respectively, then COBiassingle is computed by:
$\textrm{COBias}_{\textrm{single}}=\displaystyle\sum_{i\neq o}\bigg{|}A_{o}-A_{i}\bigg{|}$ (1)
Class accuracy $A_{i}$ is given by $\frac{1}{|\mathcal{C}_{i}|}\sum_{m\in\mathcal{C}_{i}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{C}_{i}$ is the set of indices of class $c_{i}$ examples, and $\mathbbm{1}(\boldsymbol{\cdot})$ returns 1 if the condition inside is satisfied and 0 otherwise.
However, bedsides the oddest class $c_{o}$, there could be the second oddest, the third oddest, or etc. class for a given class. Generalizing the definition, we extend the contextual oddity bias metric as:
$\textrm{COBias}={N\choose 2}^{-1}\displaystyle\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{i}-A_{j}\bigg{|}$ (2)
The contextual oddity bias metric reflects the inter-class imbalance in accuracy. This metric considers each class as the odd class to others, representing the difference between every pair of classes.
The Debiasing Problem. The problem is two-fold - we need to reduce the contextual oddity bias, while maintaining or increasing the overall test accuracy.
@@ 3 Improving LLM Outputs by Debiasing as Nonlinear Integer Programming
Figure 1: An overview of the DNIP method. We take the TREC [20] task for illustration; the ICL approach uses 1-shot prompting whose prompt is given by light yellow text boxes. The upper-left plot shows that Loc. is an odd class to the others, holding their majority wrong predictions, especially for Ent.; the upper-right plot demonstrates the stark contrast in per-class accuracy, with 100% for Loc. and 0 for Ent.. The bottom plots show that DNIP greatly reduces pairwise class accuracy differences, especially boosting accuracy for Ent. ($0\rightarrow 66\%$).
In this work, we approach the debiasing problem as a nonlinear integer programming (NIP) problem. The objective is to optimize correction weights to rectify the biased, per-class probabilities, with directly modeling the COBias metric and two accuracy-related metrics. The weights are optimized on a set of labeled examples, and are simply plugged in to inference-time probabilities. An overview of the proposed DNIP method is shown in Figure 1. Our source code will be released.
@@ 3.1 The DNIP Objective
Considering $M$ examples with known class labels, we let $\boldsymbol{p}_{m}$ denote the normalized output probability over $N$ class tokens for an example, and $\boldsymbol{p}_{m}=(p_{m1},\dots,p_{mn},\dots,p_{mN})$, $\sum_{n=1}^{N}p_{mn}=1$, $m\in\{1,\dots,M\}$. To debias $\boldsymbol{p}_{m}$, we introduce a discrete $K$-point correction scale $\boldsymbol{\omega}$ ranging from 0 to 1, to re-weight class probability. Let $\boldsymbol{\omega}=(\omega_{1},\dots,\omega_{K})$ denote the parameter for the discrete weights in the correction scale, where $\omega_{k}=k/K$. An example $10$-point scale is given by $\boldsymbol{\omega}=(0.1,\dots,1.0)$. For each class probability in $\boldsymbol{p}_{m}$, it times a correction weight chosen from $\boldsymbol{\omega}$ to obtain a debiased class probability.
@@
Instead, we model the debiasing problem as a nonlinear integer program, and search for an optimal set of class correction weights with simulated annealing. To this end, the problem boils down to selecting optimal weight indices, and we define an integer selection variable $\xi=(\xi_{1},\dots,\xi_{N})$ that selects a weight index $k$ for the $n$-th class probability in $\boldsymbol{p}_{m}$, namely:
@@
where $\beta$ and $\tau$ are parameters that control the bias and the negative PMI terms respectively, $\hat{S}^{j}$ and $S^{j}$ denote the examples of prediction $j$ and true label $j$ respectively [^1] In details, $f(\hat{S}^{j})$ is the ratio between the number of examples with prediction $j$ and the total number of examples, similarly, $f({S}^{j})$ is the ratio between the number of examples with class label $j$ and the total number of examples, $f(\hat{S}^{j},S^{j})$ is the ratio between the number of correct predictions of class $j$ and the total number of examples. Therefore, $\textrm{PMI}_{j}=\log\big{(}\frac{\mathrm{n}(\hat{S}^{j},S^{j})/N}{\mathrm{n}(\hat{S}^{j})/N\cdot\mathrm{n}(S^{j})/N}\big{)}$, where $\mathrm{n}(\cdot)$ is the count. In actual computations, we apply add-$u$ smoothing to smooth out zeros in the numerator and denominator. The value of the smoothing coefficient $u$ is selected along other parameters on a development set.. The debiased class probabilities and prediction are:
@@
The intuition is to encourage the adjusted predictions to have lower inter-class bias while improving overall accuracy. As the second term COBias has minimized the inter-class accuracy bias, the first term error rate brings the class prediction closer to the actual class, and the third term negative PMI further enforces the class prediction to be more confident when it is close to the actual class.

### v3

# Balancing Class Accuracies for Language Models in Inference Time via Nonlinear Integer Programming

@@ Abstract
Large language models are good knowledge bases but struggle to perform equally well for all classes in simple text classification tasks. This paper investigates a fundamental inference-time problem in language models: imbalanced class accuracies. We find what’s underneath the issue is a tendency to over-predict some classes while under-predicting some others, for both base LLMs and more sophisticatedly trained, larger chat LLMs. This class accuracy imbalance is difficult to solve from the root via “better” pre-training or fine-tuning strategies, but we show it can be effectively mitigated via inference-time optimization. To this end, we conceptualize and quantify the over- and under-prediction issue as the Contextual Oddity Bias (COBias), and propose the Debiasing as Nonlinear Integer Programming (DNIP) model to correct in-context learned class probabilities based on minimizing COBias and maximizing overall accuracy, without LLM parameter update. Considering that the DNIP model implicitly contains non-differentiable elements, we therefore use the simulated annealing algorithm to solve. We perform extensive evaluations on three LLMs across seven NLP classification tasks, in different prompting settings. Results show that DNIP simultaneously achieves significant COBias reduction (-27%) and accuracy improvement (+12%) over the conventional ICL approach, suggesting that inference-time mitigation of class accuracy imbalance is a promising direction to push forward LLM performances.
@@ 1 Introduction
The fact that large language models (LLMs) achieve remarkable overall test accuracy by few-shot in-context learning (ICL) [1] can obscure a large difference in individual class accuracies. Mitigating class accuracy imbalance is under-explored, but important to facilitate a smoother user experience. Not mitigating such imbalances may lead to serious damages in some cases. For example, a patient can trust an LLM about drug-drug interaction type predictions without knowing that the model over-predicts “no interaction” and has low accuracies for other categories of interactions. If the patient trusts “no interaction” when the ground truth is an actual drug interaction type, taking the drugs together can put the patient’s life at risk. Therefore, it is high time we took actions to reduce class accuracy imbalances, better without hurting overall accuracy.
It is difficult to disentangle all the factors that cause class accuracy imbalance, including biases incurred in pre-training, supervised fine-tuning, RLHF, prompting specifics, etc. Therefore, it is challenging to comprehensively resolve the root causes through pre-training or fine-tuning methods, or approaches like data selection or data augmentation. To overcome these difficulties, we conceptualize a post-hoc bias metric, the Contextual Oddity Bias (COBias; Section 2), to quantify the class accuracy imbalance based on inference-time over-predictions and under-predictions.
To reduce COBias and rectify LLM outputs, we directly operate on the class probabilities obtained from in-context learning (ICL). We propose DNIP, the Debiasing as Nonlinear Integer Programming model, to adjust class probabilities towards lower COBias and higher overall accuracy, from a combinatorial optimization viewpoint (Section 3). We formulate the debiasing problem as a discrete correction weight selection problem. In particular, DNIP re-weights class probability with an optimal set of weights selected based on both metrics of accuracy and COBias, and thus obtains much less biased predictions with higher accuracy.
On three widely used different-scale LLMs across seven NLP evaluation datasets, and in different ICL settings, we show the effectiveness of DNIP. Our key messages are as follows:
  • We introduce the COBias evaluation metric to assess the pairwise class accuracy bias in language model predictions, and reveal that LLMs of different scales and families consistently exhibit a large COBias score.
@@
  • Experiments show that DNIP achieves the best of both worlds of lower COBias and higher overall accuracy. For example, on Llama-2-13B, DNIP significantly reduces COBias (avg. $40.5\%\rightarrow 14.3\%$) and improves overall accuracy (avg. $59.4\%\rightarrow 69.2\%$) over the 1-shot ICL results, and it reduces COBias even further over the 5-shot ICL results (avg. $39.2\%\rightarrow 10.3\%$).
2 Over-prediction, Under-prediction, and the Contextual Oddity Bias
2.1 Over-prediction and Under-prediction in Classification
[…]
In the era of Large Language Models (LLMs), over-prediction and under-prediction still persist in LLMs’ predictions, though one may not simply attribute such issues to the black-box training data or the model itself. The first obstacle is that the pretraining data’s source distribution for a subgroup of data may not be accessible, making it difficult to directly minimize differences between the source distribution and a downstream target distribution. What we normally have is the pretrained model’s output distribution for downstream classes. Secondly, the process of prompting an LLM, especially when the prompt is provided with in-context learned demonstrative examples, also brings uncertainty to predictions. Ablating the potential biases in prompting is non-trivial. More inconveniently, given the relatively large model sizes, it is computationally intensive, expensive, or even infeasible to tweak the LLM training pipeline or conduct LLM fine-tuning to fit debiasing training objectives.
To see how over-predictions and under-predictions hurt LLM prediction capabilities, we show in Table 2 an example of observed over-predictions and under-predictions in Llama-2-13B. We prompt the LLM on the AGNews news classification test set in a one-shot manner, with a single demonstrative example randomly selected from the training set. Llama-2-13B greatly over-predicts Business while under-predicting Technology, lowering the LLM’s capability for accurately predicting the Tech class (Tech’s accuracy: 19%). In general, there are 2129 predictions in Business but the number of ground-truth instances in Business is 1204; there are only 246 correct Tech predictions but 1240 ground-truth Tech instances. Over-/Under-prediction happens in more downstream tasks and larger LLMs (e.g., ChatGPT), as additionally show by examples in Appendix A. These surface observations tell us how an LLM could fail in predicting certain classes, suggesting that there is still space for improvements for LLMs to predict every class with equally good accuracy. The fact that this problem is difficult to completely eliminate from the source of data, together with the inconveniences in training/fine-tuning debiased LLMs, call for a more actionable inference-time bias metric quantifying the over-/under-predictions and an accompanying debiasing method.
Pred.
World Sports Business Tech All Acc.
World 1093 64 126 3 1286 0.85
Sports 9 1247 14 0 1270 0.98
True Business 25 4 1167 8 1204 0.97
Tech 156 27 822 235 1240 0.19
All 1283 1342 2129 246 5000 -
Prec. 0.85 0.93 0.55 0.96 - -
Table 1: Over-prediction in class Business and under-prediction in class Tech in Llama-2-13B, shown by the test confusion matrix for AGNews.
Pred.
World Sports Business Tech All Acc.
World 1100 50 67 69 1286 0.86
Sports 16 1240 11 3 1270 0.98
True Business 44 2 1028 130 1204 0.85
Tech 55 7 138 1040 1240 0.84
All 1215 1299 1244 1242 5000 -
Prec. 0.91 0.95 0.83 0.84 - -
Table 2: Significantly reduced over-predictions in class Business and improved test accuracy of the under-predicted class Tech for AGNews after DNIP.
2.2 Conceptualizing Over-prediction and Under-prediction into the Contextual Oddity Bias
In Table 2, we observe that Llama-2-13B greatly over-predicts Business while under-predicting Technology (Tech). Underneath the surface over-/under-prediction issue, we can see the most significant contributor is the 822 ground-truth instances of Technology, which are wrongly predicted as Business, leading to the model’s low accuracy (19%) in Technology [^1] It simultaneously leads to a fairly low precision score of 55% in Business. In fact, the low precision score in Business and the low recall/accuracy score in Technology co-reflect each other. In other words, class accuracies are essential indicators for class performances. All current evaluation metrics are computed by Python scikit-learn. Note that by scikit-learn, class recall is the same as class accuracy.. On that note, we are able to precisely describe the surface over-prediction and under-prediction issue, by decoupling the problem into which output class that a ground-truth class is most wrongly predicted as, suggesting the model’s inference bias towards an output class given a ground-truth class. In the AGNews example, Business is what ground-truth Technology is most wrongly predicted as, suggesting the model’s inference bias to predict Technology as Business.
To quantify this inference bias, we first conceptualize odd classes. An odd class is an output class at inference time which holds the most wrong predictions of a ground-truth class $c_{i}$. That is, an odd class is always relative to a ground-truth class. For $N$ classes and $M$ test instances, let $y$ and $\hat{y}$ denote the ground-truth class and the prediction respectively, and let $u_{c_{i},c_{j}}$ denote the number of ground-truth $c_{i}$ instances that are wrongly predicted as $c_{j}$,
$u_{c_{i},c_{j}}=\displaystyle\sum\nolimits_{m}\mathbbm{1}\{y_{m}=c_{i},\hat{y}_{m}=c_{j}\},$ (1)
$odd_class|(c_{i}):=\operatorname*{arg\,max}_{c_{j}}u_{c_{i},c_{j}},$ (2)
[…]
For this goal, it could be direct to use the rate of wrong predictions in an odd class as a measurement of its “hurt” to the accuracy of a ground-truth class, but minimizing these rates alone may not adequately mitigating class accuracy imbalances, which is an ultimate goal of reducing over-/under-predictions (More analyses on this measurement are in Appendix B.). A simpler yet better measurement is the absolute accuracy difference between the odd class and the ground-truth class. For example, knowing that Business is odd to Tech, the 78% accuracy difference between Business (97%) and Tech (19%) can serve as a measurement of how much Tech is under-predicted. The 19% accuracy of Tech also implicitly reflects the many predictions in its odd class Business. It should be noted that measuring these accuracy differences sometimes fails to capture the hidden ground-truth-and-odd relationships. For instance, World is the odd class to ground-truth Business, but the -12% accuracy difference can not utter this fact. Despite this drawback, directly minimizing the accuracy differences will in return significantly reduce predictions in odd classes and improve under-predicted classes’ accuracies.
@@ 2.2 Conceptualizing Over-prediction and Under-prediction into the Contextual Oddity Bias
The Contextual Oddity Bias. To this end, we formulate the definition of Contextual Oddity Bias (COBias). First, the absolute value of the accuracy difference between a ground-truth class and its odd class is termed as COBiassingle. COBiassingle measures a model’s tendency to over-predict an answer that is caught as odd than the correct answer (or equivalently, a tendency to under-predict the correct answer). Notably, for LLMs, an odd answer may be a manifest of surface patterns from prompts or biases in the pretraining data. In addition, the odd classes in a same downstream task can vary across different models. That’s why we use “contextual” to mean that the oddity of an answer depends on the model or task.
@@
Formally, given an $N$ classes and $M$ instances, let $x_{m}$ denote an instance whose label is $y_{m}$. By few-shot prompting, the model’s output token probability over $N$ classes $c_{1},\dots,c_{N}$ is $\boldsymbol{p}_{m}=(p_{m1},\dots,p_{mN})$. $\boldsymbol{p}_{m}$ is normalized over $N$ classes, $\sum_{n=1}^{N}p_{mn}=1$, $m\in\{1,\dots,M\}$. Then the prediction $\hat{y}_{m}$ is $\operatorname*{arg\,max}_{j\in\{1,\dots,N\}}\boldsymbol{p}_{m}$. For a given class $c_{i}$, let $A_{\textrm{odd}_{i}}$ and $A_{c_{i}}$ denote the accuracies for the odd class $\textrm{odd}_{i}$ and $c_{i}$ respectively, the COBiassingle metric is computed as:
$\textrm{odd}_{i}=\verb|odd_class|(c_{i}),$ (3)
$\textrm{COBias}_{\textrm{single}}=\frac{1}{N}\displaystyle\sum\nolimits_{i}\bigg{|}A_{\textrm{odd}_{i}}-A_{c_{i}}\bigg{|},$ (4)
where class accuracy $A_{c_{i}}=\frac{1}{|\mathcal{S}_{c_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{c_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{S}_{c_{i}}$ is the set of indices for class $c_{i}$ instances. Similarly, $A_{\textrm{odd}_{i}}=\frac{1}{|\mathcal{S}_{\textrm{odd}_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{\textrm{odd}_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{S}_{\textrm{odd}_{i}}$ is the set of indices for class $\textrm{odd}_{i}$ instances.
Moreover, besides the odd class, which is the most-frequent wrongly predicted class, there could also be a secondly most-, or third most-, to the least-frequent wrongly predicted class. For a ground-truth class, there are the odd, the second odd, to the least odd classes. For example, for ground-truth Tech (1240 instances), Business is the odd (822), while World is the second odd (156), and Sports is the least odd (27). Then, we extend the accuracy difference between the single odd class and a ground-truth class to every pair of classes, and absolute value can be taken. This arrives at the ultimate the COBias metric as follows.
$\textrm{COBias}={N\choose 2}^{-1}\displaystyle\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{c_{i}}-A_{c_{j}}\bigg{|}.$ (5)
This COBias metric directly reflects the inter-class imbalance in accuracy. Indeed, the absolute value of accuracy differences of some pairs in Equation 5 may mask the most frequent wrongly predicted class (the odd class). Nonetheless, it reflects a well-rounded measurement of all pairs of classes, whether involving odd classes or not, so it is a powerful metric when we strongly demand minimal class accuracy imbalance.
In a nutshell, the debiasing problem is minimizing COBias (Eq. 5) while maintaining or increasing the overall inference accuracy. This post-hoc bias evaluation metric will enable easy-to-operate inference-time/post-hoc output likelihood optimization, lifting the heavy computations required for training-wise data selection or model parameter updates.
@@ 3 Improving LLM Outputs by Debiasing as Nonlinear Integer Programming
Figure 1: An overview of the DNIP method. We take the TREC [5] task for illustration; the ICL approach uses 1-shot prompting whose prompt is given by light yellow text boxes. The upper-left plot shows that Loc. is an odd class to the others, holding their majority wrong predictions, especially for Ent.; the upper-right plot demonstrates the stark contrast in class accuracy, with 100% for Loc. and 0 for Ent.. The bottom plots show that DNIP greatly reduces pairwise class accuracy differences, especially boosting accuracy for Ent. ($0\rightarrow 66\%$).
In this work, we approach the debiasing problem as a nonlinear integer programming (NIP) problem. The objective is to optimize correction weights to rectify the biased, class probabilities, by directly minimizing the COBias metric while maximizing two accuracy-related metrics. The weights are optimized on a set of labeled examples, and are simply plugged in during test time. An overview of the proposed DNIP method is shown in Figure 1.
@@ 3.1 The DNIP Objective
To debias $\boldsymbol{p}_{m}$, we introduce a discrete $K$-point correction scale $\boldsymbol{\omega}$ ranging from 0 to 1, to re-weight class probability. Let $\boldsymbol{\omega}=(\omega_{1},\dots,\omega_{K})$ denote the parameter for the discrete weights in the correction scale, where $\omega_{k}=k/K$. An example $10$-point scale is given by $\boldsymbol{\omega}=(0.1,\dots,1.0)$. For each class probability in $\boldsymbol{p}_{m}$, it multiplies by a correction weight chosen from $\boldsymbol{\omega}$ to obtain a debiased class probability.
@@
Instead, we model the debiasing problem as a nonlinear integer program, and search for an optimal set of class correction weights with simulated annealing. Without loss of generality, we boil down the problem to selecting optimal weight indices towards minimizing the pairwise class accuracy imbalance, while maximizing the overall accuracy and the pointwise mutual information between class labels and predictions. Formally, we define an integer selection variable $\xi=(\xi_{1},\dots,\xi_{N})$ that selects a weight index $k$ for the $n$-th class probability in $\boldsymbol{p}_{m}$, namely:
@@
where $\beta$ and $\tau$ are parameters that control the bias and the negative PMI terms respectively, $\hat{S}^{j}$ and $S^{j}$ denote the examples of prediction $j$ and true label $j$ respectively. In details, $f(\hat{S}^{j})$ is the ratio between the number of examples with prediction $j$ and the total number of examples, similarly, $f({S}^{j})$ is the ratio between the number of examples with class label $j$ and the total number of examples, $f(\hat{S}^{j},S^{j})$ is the ratio between the number of correct predictions of class $j$ and the total number of examples. Therefore, $\textrm{PMI}_{j}=\log\big{(}\frac{\mathrm{n}(\hat{S}^{j},S^{j})/N}{\mathrm{n}(\hat{S}^{j})/N\cdot\mathrm{n}(S^{j})/N}\big{)}$, where $\mathrm{n}(\cdot)$ is the count. In actual computations, we apply add-$u$ smoothing to smooth out zeros in the numerator and denominator. The value of the smoothing coefficient $u$ is selected along other parameters on a development set.
The debiased class probabilities and prediction are:
@@
The intuition is to encourage the adjusted predictions to have lower inter-class bias while improving overall accuracy. As the second term COBias has minimized the inter-class accuracy bias, the first term error rate brings the class prediction closer to the actual class, and the third term negative PMI further enforces the class prediction to be more confident when it is close to the actual class. The PMI term can also be seen as a constraint to penalize each individual class.